OpenStudy (anonymous):

A box contains 20 cards numbered 1-20. You select a card. Without replacing the first card, you select a second card. Find each probability P(1, then 20) P(even, then 7)

OpenStudy (raden):

P(1) = 1/20 and P(20) = 1/19 therefore P(1 and 20) = P(1) * P(20|1) = 1/20 * 1/19 = .... part b is samething like part a

OpenStudy (anonymous):

So the answer would be 1/380

OpenStudy (raden):

yes, for the 1st one

OpenStudy (anonymous):

Ok thank you, would you mind helping me set up the second one please

OpenStudy (raden):

1-20, how many even numbers are there ?

OpenStudy (anonymous):

10

OpenStudy (raden):

yup, so P(even) = 10/20 and P(7) = 1/19, right ?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Then multiply them together ?

OpenStudy (raden):

yup

OpenStudy (anonymous):

10/380

OpenStudy (raden):

you got it!

OpenStudy (anonymous):

Thanks for your help!!

OpenStudy (raden):

yoyu're welcome